Tasty Bakery sells three kinds of muffins: chocolate chip muffins at 30 cents each, oatmeal muffins at 35 cents each, and cranberry muffins at 40 cents each. Charles buys some of each kind and chooses twice as many cranberry muffins as chocolate chip muffins. If he spends $4.65 on 13 muffins, how many oatmeal muffins did he buy?
Answer by jorel1380(2518)

cc+t+cr=13
30cc+35t+40cr=465
cr=2cc
cc+t+2cc=13
3cc+t=13
30cc+35t+80cc=465
110cc+35t=465
105cc+35t=455
5cc=10
cc=2
cr=4
t=7
Charles bought 2 chocolate chip muffins, 4 cranberry, and 7 oatmeal..